The area of a rectangle is 44 yards (squared) , and the length of the rectangle is 3 yd less than twice the width. Find the dimensions of the rectangle.
Answer by swincher4391(1107)

A = L*W
L = 2W - 3
A = (2W-3)*W
44 = 2W^2 - 3W
2W^2 - 3W - 44 = 0
((3 +- sqrt((-3)^2 - 4*2*-44))/4

(3 +- sqrt( 9 + 352))/4
(3 +- 19)/4
W= -16/4 and 22/4
However, W must be positive as we are talking about a measurement.
So W = 22/4 = 5.5
Plug that back into our original 44 = L * 5.5
44/5.5 = 8
So our dimensions are 8 x 5.5.